Summary of TED runs on February 18th/19th 2010

Run Number BCID # shots TT HV TT Timing IT Timing Miscellaneous
66183 3156 28 350 Nominal Nominal High intensity
66184 3156 26 350 -3ns -3ns lower intensity
66185 3156 20 350 +3ns Nominal  
66186 3156 20 350 +6ns +3ns  
66187 3156 3 350 +9ns +3ns  
66188 3156 31 350 +9ns +3ns new run after reinstallation of firmware on ittell09, not clear which delays were loaded
66189 3156 14 350 +9ns +3ns  
66190 3156 30 350 +9ns +3ns  
66191 3156 1 350 +9ns +3ns  
66192 3156 23 350 +9ns +3ns  
66193 3156 17 350 +9ns +3ns  
66194 3156 22 150 +9ns +3ns  
66195 3156 15 150 +9ns +3ns  
66196 3156 23 150 +6ns +3ns  
66197 3156 16 150 +3ns -+3ns reset of ITTELL09 after bad magic pattern errors
66198 902? 4? 150 -3ns +3ns looks like there were no shots
66199 3156 74 150 Nominal Nominal back on beam 2 after 2 hours

-- MarkTobin - 19-Feb-2010

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2010-03-01 - MarkTobin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback